OpenStudy (anonymous):

Show all of your steps and solve 0.25r - 0.125 + 0.5r = 0.5 + r @texaschic101

OpenStudy (anonymous):

Make sure you show all of the steps for my benefit

OpenStudy (skullpatrol):

For your maximum benefit you should TRY to do it yourself.

OpenStudy (texaschic101):

first combine like terms

OpenStudy (anonymous):

ok hold on

OpenStudy (texaschic101):

just on the left side of the equal sign, there are no like terms on the right side

OpenStudy (texaschic101):

do you know what like terms are ?

OpenStudy (anonymous):

0.25r - 0.125 + 0.5r = 0.5 + r -0.5r -0.5r 0.20r - 0.125 = 0.5 + r

OpenStudy (texaschic101):

let me show you.... .25r - .125 + .5r = .5 + r on the right side of the equal sign, add your r terms. (.25r + .5r)

OpenStudy (texaschic101):

it should look like this after adding your r terms... .30r - .125 = .5 + r do you understand ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

now you want to get your r terms on one side and everything else on the other side. So you want to move the r term on the right side to the left side. Do you know how to do that ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

ok.....to do that, you do the opposite of what the r says. I know that is confusing. I will try to explain. The r on the right side says + r so do the opposite and -r on both sides. This will cancel out the r on the right side. So subtract r from both sides

OpenStudy (texaschic101):

it should look like this.... .30r - .125 - r = .5 + r - r as you can see the r on the right side cancels out Does that make sense to you ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

now to make things easier combine your like terms

OpenStudy (texaschic101):

but remember, the r is 1r. So when combining like terms on the right side, remember this

OpenStudy (texaschic101):

am I confusing you ? I am not the best at explaining things

OpenStudy (anonymous):

yes lol Im kind of confused it would be easier if you stacked your steps it would make it easier on me Step 1 step 2 step3

OpenStudy (texaschic101):

.30r - .125 - r = .5 combine like terms do you know what .30r - 1r equals ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

ok .30r - 1r subtract .30 from 1 and put a negative sign in front

OpenStudy (anonymous):

.70

OpenStudy (texaschic101):

you forgot the negative but good job...you just combined like terms... so now we have : - .70r - .125 = .5 have I lost you, or do you understand ?

OpenStudy (anonymous):

im understanding but at the end will you work it out without explaining it helps

OpenStudy (texaschic101):

at the end I will. now we want the r on one side and everything else on the other side. So we have to move the .125 to the other side. So we will do the opposite. It says -.125 so we have to +.125 (add .125 to both sides). Can you do that ?

OpenStudy (anonymous):

.625

OpenStudy (texaschic101):

-.70r -.125 + .125 = .5 + .125......it is not 5, it is .5 do not forget your decimal. Combine like terms.

OpenStudy (texaschic101):

remember the .125's cancel each other out on the right side.

OpenStudy (anonymous):

I didn't forget the decimal try adding .125+.5 on a calculator it adds up to .625

OpenStudy (texaschic101):

you right...my mistake..... so what is your equation once you combine like terms

OpenStudy (anonymous):

-70r = .625

OpenStudy (texaschic101):

correct. You are doing good. Now we want r by itself. It is saying multiply, so we do the opposite and divide. -.70r/-.70 = .625/-.70 can you do this ?

OpenStudy (anonymous):

.89

OpenStudy (texaschic101):

the -.70's on the right side, when divided = 1, so your left with 1r or just r. so you have r = - .892 or -.89, it depends on how many decimal places you want to go to. Remember a positive divided by a negative equals a negative.

OpenStudy (texaschic101):

I will write the problem out for you now, give me a second

OpenStudy (anonymous):

I think we messed up when we done .25+.5

OpenStudy (anonymous):

gotta go walk my dog ill be back

OpenStudy (texaschic101):

.25r - .125 + .5r = .5 + r .75r - .125 = .5 + r .75r - r = .5 + .125 -.25r = .625 r = .625/-.25 r = - 2.5

OpenStudy (texaschic101):

let me check it.... .25r - .125 + .5r = .5 + r .25(-2.5) - .125 + .5(-2.5) = .5 + (-2.5) -.625 - .125 - 1.25 = .5 - 2.5 -2 = -2 (correct) r = -2.5 sorry about the mistake earlier.